OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos 2 (theta) = 2 cos^2 (theta) - 1

zepdrix (zepdrix):

Rewrite it like this. \[\cos(2\theta)=\cos(\theta + \theta)\] Then you need to remember your angle sum identity for cosine :) \[\cos(\alpha+\beta) = \cos \alpha \cos \beta - \sin \alpha \sin \beta\] See if you can figure it out from there! :)

OpenStudy (anonymous):

So, 2+2? for a + b?

zepdrix (zepdrix):

Theta + Theta for a and b :\